Solve the system of equations: {4x+y+5z=6}
{2x+2y-z=11}
{x+2y+2z=6}

Answer by Edwin McCravy(20056)

Solve the system of equations:
{4x+y+5z=6}
{2x+2y-z=11}
{x+2y+2z=6}
There are several different ways to solve this.
So I can only give you the answer. If you want
to see the solution, post again telling which method
you are studying, addition and substitution, Gaussian
elimination matrix method, inverse matrix method,
or Cramer's rule.
The answer, by any method, is
(x,y,z) = (2,3,-1)
Edwin